### v2

# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

Abstract
The rapid growth of the low-altitude economy has driven the widespread adoption of unmanned aerial vehicles (UAVs). This growing deployment presents new challenges for UAV trajectory planning in complex urban environments. However, existing studies often overlook key factors, such as urban airspace constraints and economic efficiency, which are essential in low-altitude economy contexts. Deep reinforcement learning (DRL) is regarded as a promising solution to these issues, while its practical adoption remains limited by low learning efficiency. To overcome this limitation, we propose a novel UAV trajectory planning algorithm that integrates DRL with the large language model (LLM) reasoning to enable safe, compliant, and economically viable trajectory planning. Specifically, we model the data collection task as a partially observable Markov decision process, explicitly incorporating obstacle avoidance, regulation awareness, and energy constraints. We design a hybrid optimization algorithm based on the soft actor-critic algorithm and LLM reasoning to enable adaptive decision-making in uncertain and dynamic environments. Experimental results demonstrate that our algorithm achieves the best overall performance, with the highest data collection rate (99.50%), almost zero collision avoidance rate and regulation violation rate, perfect successful landing rate (100%), and the lowest energy consumption rate (76.95%). These results validate the effectiveness of our algorithm in addressing UAV trajectory planning key challenges under constraints of the low-altitude economy networking.
Index Terms:
Data Collection, Large Language Model, Low-Altitude Economy, Reinforcement Learning, Trajectory Planning
1 Introduction
With the rapid advancement of the low-altitude economy, low-altitude airspace has emerged as a critical strategic resource to drive urban intelligence, industrial upgrading, and digital economic transformation [1]. This evolving domain integrates diverse sectors such as general aviation and urban air mobility, wherein unmanned aerial vehicles (UAVs) play an increasingly vital role in applications including logistics, agricultural monitoring, infrastructure inspection, and environmental protection [2]. Efficient utilization of this low-altitude airspace presents unprecedented opportunities for industry growth and technological innovation.
In this context, UAV-based data acquisition represents a fundamental technology enabling high-precision, multi-modal environmental sensing [3]. The UAV data acquisition process typically involves task assignment, flight trajectory planning, sensor data collection, and data transmission and processing. Effective trajectory planning is crucial to maximize area coverage and data quality, thereby supporting downstream applications with accurate and timely information [4].
Despite its potential, trajectory planning for UAV data acquisition in low-altitude airspace faces the following important challenges:
Challenge 1: Robust Obstacle Avoidance in Urban Airspaces. Due to the existence of airspace reuse in the low-altitude economy [36], operating in low-altitude urban environments exposes the data collection UAVs (DCU) not only to static obstacles (e.g., buildings) but also to dynamic obstacles (e.g., other UAVs). These obstacles are often partially observable and highly uncertain. Moreover, the DCU must also avoid restricted no-fly zones (NFZs) over sensitive areas. Ensuring robust obstacle avoidance under such conditions remains a core challenge. Traditional optimization methods typically lack the adaptability to handle such stochastic environments, and standard deep reinforcement learning (DRL) struggles to accumulate sufficient successful trajectories, leading to low learning efficiency [5].
Challenge 2: Intelligent Planning for the Economy with Compliance. In the low-altitude economy, the DCU must not only maximize data acquisition but also minimize energy consumption to ensure cost-effectiveness. Moreover, the DCU must adhere to evolving, location-specific regulations where the DCU is to reduce speed to minimize noise [6]. These decisions introduce additional requirements on mission efficiency and regulatory compliance. Traditional methods struggle with multi-objective adaptation under uncertainty, while conventional DRL approaches often fail to incorporate domain knowledge needed to make economically viable decisions.
Challenge 3: Real-Time Decision-Making Under Uncertainty. Trajectory planning in real-world scenarios requires real-time decision making across multiple dimensions under dynamic and uncertain conditions, including obstacle avoidance, route selection, and data collection. Traditional optimization methods are often computationally intensive and lack responsiveness, whereas standard DRL policies may require extensive exploration due to task complexity.
Research Gap. Most existing works have focused on specific aspects of UAV trajectory planning while neglecting others. In terms of traditional algorithms, [7] and [9]-[13] optimized communication and energy efficiency but did not consider obstacle avoidance. Although [8] and [14] incorporated obstacle avoidance, they were limited to static obstacles and overlooked dynamic threats like other UAVs. Meanwhile, DRL has been increasingly adopted for UAV-related optimization tasks due to its capacity to adaptively control in dynamic environments [18]. Studies using DRL, such as [27], focused on static obstacle avoidance or energy efficiency, yet lacked mechanisms for dynamic obstacle avoidance. Besides, [19]-[26] neglected critical aspects such as obstacle avoidance and adherence to regulatory constraints, limiting their applicability in real-world low-altitude environments. In parallel, large language models (LLMs) have also been introduced in mobile and aerial network scenarios to enhance intelligent decision-making and task coordination  [15]-[17]. Although some existing studies  [28]-[32] applied the LLM to UAV trajectory planning, they primarily focused on obstacle avoidance while neglecting compliance requirements and the need for data collection.
In summary, existing methods either focus on isolated objectives or assume simplified environments, and thus fail to address all the challenges mentioned above of the DCU trajectory planning. To address these limitations, this paper proposes a novel hybrid DRL and the LLM algorithm for safe and economical DCU trajectory planning in low-altitude airspace with compliance-awareness. Specifically, we introduce an LLM to enhance the DCU’s decision-making capability in complex, uncertain scenarios where rule compliance, contextual awareness, and obstacle interpretation are critical. While the LLM offers strong reasoning abilities and can infer safe actions based on high-level semantic input, it lacks the ability to learn optimal control strategies through interaction with dynamic environments [33]. Conversely, DRL excels at learning adaptive control policies via trial and error, but struggles to incorporate abstract constraints such as regulatory rules [34]. Therefore, we combine DRL with LLM to exploit their complementary strengths. The main contributions are as follows:
  • Mathematical modeling of DCU trajectory planning problem for low-altitude data collection. We model the DCU trajectory planning for low-altitude data acquisition task as a mathematical problem. This problem captures the robust obstacle avoidance requirement under both static and dynamic obstacles mentioned in Challenge 1.
  • Problem formulation as a partially observable Markov decision process (POMDP). We formulate the mathematical problem as a POMDP to account for partial observability in the low-altitude airspace. To capture specific requirements, such as the robust obstacle avoidance mentioned in Challenge 1 and the requirement of economy and compliance mentioned in Challenge 2, we carefully design the state space, action space, and reward function.
  • A hybrid SAC and LLM-based trajectory planning algorithm. We develop a novel optimization algorithm that combines the soft actor-critic (SAC) algorithm with an LLM. By combining the adaptive decision-making capabilities of SAC with the contextual reasoning strength of the LLM, the proposed algorithm enables the DCU to perform trajectory planning in real time, with robust obstacle avoidance, regulatory compliance, and economical data collection in uncertainty, thus addressing Challenge 3.
Extensive experiments demonstrate that our proposed algorithm outperforms baseline approaches. Compared to the best-performing baseline, our algorithm improves the data collection rate by 2.9%, reduces the collision rate and regulation violation rate to almost 0 %, maintains 100% success landing rate, and lowers the energy consumption rate by 1.9%, thereby effectively addressing the above three challenges.
The remainder of this paper is organized as follows. Section 2 reviews the related work. Section 3 describes the overall system architecture, while Section 4 presents the formal problem definition. Section 5 introduces the POMDP-based modeling approach and elaborates on the proposed algorithm. Section 6 reports the experimental results and performance evaluations. Finally, Section 7 concludes the paper and outlines future research directions.
2 Related Work
In this section, we provide related works on planning UAV trajectories using traditional algorithms, DRL, and the LLM. The summary of related work is shown in Table I.
2.1 UAV Trajectory Planning with Traditional Algorithms
Qin et al. [7] employed a joint optimization algorithm using Dinkelbach’s method and block coordinate descent (BCD) to optimize trajectory, power, bit allocation, and phase shift in reconfigurable intelligent surface (RIS)-assisted UAV-enabled mobile edge computing (MEC) systems. Pan et al. [8] proposed a joint UAV power and trajectory planning framework using an improved non-dominated sorting genetic algorithm-II with K-means initialization operator and variable dimension mechanism (NSGA-II-KV) and an improved particle swarm optimization with normal distribution initialization, genetic mechanism, differential mechanism, and pursuit operator (PSO-NGDP). Pervez et al. [9] studied a multi-UAV MEC system and applied game theory and successive convex approximation (SCA) for joint optimization. Zhang et al. [10] used an elite-driven differential evolution algorithm with dynamic population evolution to jointly optimize UAV deployment and flight planning in Internet of Things (IoT) data collection. Heo et al. [11] incorporated an NFZ avoidance constraint into UAV trajectory planning and used quadratic transform and SCA for joint optimization. Li et al. [12] applied a swarm intelligence-based algorithm (SIA) to a multi-objective problem involving collaborative beamforming for energy-efficient data harvesting. Fu et al. [13] adopted the distributed approximate Newton (DANE) method and BCD to optimize UAV trajectory, device scheduling, and resource allocation. Wang et al. [14] utilized an ant colony optimization (ACO)-based algorithm for global obstacle avoidance and communication in RIS-assisted UAV systems.
2.2 UAV Trajectory Planning with DRL
Silvirianti et al. [19] proposed a layerwise quantum-based DRL (LQ-DRL) method for UAV trajectory planning, user grouping, and power allocation. Chen et al. [20] formulated a multi-stage mixed-integer nonlinear programming problem and proposed the joint trajectory planning and resource allocation by DRL for UAV-assisted MEC. Ning et al. [21] addressed distributed trajectory control using DRL and game theory in multi-UAV MEC networks. Song et al. [22] modeled a multi-objective Markov decision process with a trace-based experience replay scheme for energy-efficient UAV trajectory planning. Ding et al. [23] proposed multi-agent advantage actor-critic with deep deterministic policy gradient (DDPG) and deep attention-based inference (DAI) schemes for spectrum allocation and UAV trajectory design. Wang et al [24] developed a safe DRL algorithm based on a constrained Markov decision process for UAV trajectory planning. He et al. [25] proposed a distributed multi-objective trajectory design method using distributed DRL for trajectory planning of multi-UAV. Liu et al. [26] applied multi-agent RL (MARL) and SCA to optimize UAV trajectory in a cell-free space-air-ground integrated network. Ning et al. [27] used constrained DRL to optimize UAV movement and connection assignment in wireless-powered IoT networks.
2.3 UAV Trajectory Planning with LLM
Zhong et al. [28] presented a vision-based autonomous planning system for UAVs in dynamic environments. It combined lightweight object detection and Kalman filtering to track moving obstacles. Then they explored integrating an LLM for more natural human-UAV interaction. Phadke et al. [29] explored the integration of the LLM into UAV control, enabling natural language interaction for more intuitive and accessible operations. It highlighted potential benefits in decision-making, mission planning, and safety, and presented a development framework with proof-of-concept demonstrations using existing LLMs and simulation platforms. Xiao et al. [30] introduced a foundation model-guided trajectory planning framework that integrated LLMs and vision-language models (VLMs) for enhanced semantic reasoning and visual perception in UAV navigation. Samma et al [31] proposed a novel LLM-based visual trajectory planning method for UAVs operating in indoor GPS-denied environments. By interpreting camera data on detected persons and depth, the fine-tuned LLM generated safe and efficient flight paths. Cai et al. [32] presented a hybrid autonomous UAV landing system that combined LLM with model predictive control (MPC) to improve safety and precision in dynamic, unstructured environments.
2.4 Discussion
Although numerous studies have explored UAV trajectory planning for applications such as data collection, obstacle avoidance remains an under-addressed challenge. For instance, while works [7]-[10] focused on joint optimization of resources and trajectories, they either entirely overlooked obstacle avoidance or considered only static obstacles, failing to avoid dynamic obstacles in the low-altitude airspace. Although [11] incorporated NFZ, it neglected other essential regulatory constraints. Similarly, [12]-[14] did not integrate any obstacle avoidance mechanisms, or again, only addressed static obstacles. Studies [19]-[23] ignored obstacle avoidance altogether, and although [24] introduced learning-based safety mechanisms, it did not jointly consider energy constraints and obstacle avoidance. Furthermore, [25] and [26] omitted obstacle considerations, while [27] did not account for dynamic obstacles such as other UAVs. Additionally, although studies [28]-[30] employed an LLM for UAV trajectory planning, they only considered the avoidance of a single obstacle in the planning process. While [31] and [32] offered a more comprehensive approach to obstacle avoidance, they did not account for compliance, low resource consumption, or data collection requirements.
To overcome these issues, this paper introduces an innovative algorithm that integrates DRL with the LLM to enable safe, compliance-aware, and cost-effective UAV trajectory planning in low-altitude airspace.
TABLE I: Comparison of Related Works on UAV Trajectory Planning
Ref. Main Method / Algorithm Obstacle Avoidance Compliance
Traditional
Qin et al. 2023 [7] Dinkelbach’s method, BCD
Pan et al. 2024 [8] NSGA-II-KV, PSO-NGDP Static only
Pervez et al. 2024 [9] Game theory, SCA
Zhang et al. 2024 [10] Evolutionary algorithm
Heo et al. 2024 [11] Quadratic transform, SCA
Li et al. 2024 [12] SIA
Fu et al. 2025 [13] DANE, BCD
Wang et al. 2025 [14] ACO Static only
DRL
Silvirianti et al. 2024 [19] LQ-DRL
Chen et al. 2024 [20] DRL
Ning et al. 2024 [21] DRL, game theory
Song et al. 2024 [22] Multi-objective RL
Ding et al. 2024 [23] DDPG, DAI
Wang et al. 2024 [24] Safe DRL
He et al. 2024 [25] DRL
Liu et al. 2025 [26] MARL, SCA
Ning et al. 2025 [27] Constrained DRL Static only
LLM
Zhong et al. 2024 [28] Apply LLM directly Dynamic only
Phadke et al. 2024 [29] Apply LLM directly Static only
Xiao et al. 2025 [30] LLM, VLM Static only
Samma et al. 2025 [31] Fine-tuned LLM Both Dynamic and static
Cai et al. 2025 [32] LLM, MPC Both Dynamic and static
  • Note: ✗ indicates not considered.
Figure 1: System model of low-altitude UAV data collection. The DCU collects data from GE while avoiding static obstacles (BZs, NFZs, RZs) and dynamic obstacles (OUs) under energy and compliance constraints.
TABLE II: Notations and Descriptions
Notation Description
$\mathbb{A}=(X,Y,Z)$ Area where the DCU conducts data collection, with length $X$, width $Y$, and height $Z$
$\mathbb{T}\mathbb{A}=({{p}_{\mathbb{T}\mathbb{A}}},{{l}_{\mathbb{T}\mathbb{A}}},{{w}_{\mathbb{T}\mathbb{A}}})$ Take-off area location with lower-left coordinate point, length ${l}_{\mathbb{T}\mathbb{A}}$}, and width ${w}_{\mathbb{T}\mathbb{A}}$}
$\mathbb{L}\mathbb{A}=({{p}_{\mathbb{L}\mathbb{A}}},{{l}_{\mathbb{L}\mathbb{A}}},{{w}_{\mathbb{L}\mathbb{A}}})$ Landing area location of the DCU
${{\mathbb{R}}_{XX,x}}=({{p}_{XX,x}},$ ${{l}_{XX,x}},{{w}_{XX,x}})$ Location of the $x$-th XX in $\mathbb{A}$ (XX = NFZ, BZ, or RZ)
${N}_{XX}$} Number of XX (XX = GE, OU, NFZ, BZ, or RZ)
${{L}_{GE,i}}=({{x}_{GE,i}},$ ${{y}_{GE,i}},0)$ Location of the $i$-th GE
$D{{V}_{G{{E}_{i}},t}}$ Data volume of the $i$-th GE at time $t$
$T{{P}_{GE,i}}$ Transmit power of the $i$-th GE
${{L}_{XX,t}}=({{x}_{XX,t}},$ ${{y}_{XX,t}},{{z}_{XX,t}})$ Location of XX at time $t$ (XX = DCU or OU)
$SL=({{x}_{SL}},{{y}_{SL}},{{z}_{SL}})$ Starting location of the DCU
$EL=({{x}_{EL}},{{y}_{EL}},{{z}_{EL}})$ Ending location of the DCU
$H$ Fixed flight altitude of the DCU and OUs
PRPR Perception radius of the DCU
${v}_{\max}$} Maximum flight speed of the DCU and OUs
${{v}_{XX}}(t)=({{v}_{XX}}{{(t)}_{x}},$ ${{v}_{XX}}{{(t)}_{y}},{{v}_{XX}}{{(t)}_{z}})$ Flight speed of XX at time $t$ (XX = DCU or OU)
${v}_{\text{limit}}$} Speed limit in ${\mathbb{R}}_{RZ}$} for the DCU
${E}_{total}$} Total energy the DCU has
${E}_{\text{limit}}$} Maximum energy that the DCU can consume
$E(t)$ Energy consumed during time slot $\Delta t$
$e(t)$ Remaining energy at time $t$
${{e}_{fly}}(t)$ Flight energy consumption during $\Delta t$ starting at $t$
${{e}_{hover}}(t)$ Hover energy consumption during $\Delta t$ starting at $t$
$\Delta t$ Fixed time slot duration
$T$ Total time of the data collection
${t}_{fly}$} Time used for flight in $\Delta t$
${t}_{col}$} Time used for data collection in $\Delta t$
$T{{D}_{u,G{{E}_{i}}}}(t)$ Data transmitted between DCU and GEi during $\Delta t$ starting at $t$
${{B}_{u,G{{E}_{i}}}}(t)$ Effective rate between DCU and GEi during $\Delta t$
$\left\|CP1-CP2/A\right\|$ Distance between two coordinate points or to area AA
$\wedge$ Logical AND
$\odot$ Hadamard product
3 System Description
In this section, we provide a comprehensive description of the system, including the system model and the associated models for UAV mobility, data transmission, and energy consumption. The system model defines the main entities involved. The models for mobility, data transmission, and energy consumption characterize the UAV’s movement, communication behavior, and energy usage, respectively, in a manner consistent with real-world operations. This modeling approach enhances the credibility of experimental results presented in Section 6. The notation used throughout the remainder of the paper is summarized in Table II.
3.1 System Model
The system model is shown in Fig. 1. The DCU collects data in area $\mathbb{A}$. The length, width, and height of $\mathbb{A}$ are $X$, $Y$, and $Z$, respectively. The system entities include the data collection UAV (DCU), other UAVs (OUs), and ground equipment (GE), which are detailed as follows.
Data Collection UAV (DCU). The DCU is responsible for the data collection task in the low-altitude economy and flies at a fixed altitude $H$. It takes off from $\mathbb{T}\mathbb{A}$, then conducts data collection in $\mathbb{A}$, and eventually reaches $\mathbb{L}\mathbb{A}$. When conducting data collection, the DCU is not allowed to fly into NFZs. Furthermore, when the DCU flies into residential zones (RZs), to ensure noise control, its flight speed must be less than $v_{\text{limit}}$} [28]. The DCU must also ensure that it does not collide with building zones (BZs) and other UAVs (OUs) in the low-altitude economic networking. Therefore, we assume that the DCU is equipped with a sensing device with a perception radius of $PR$. Through this device, the DCU can detect the positions and flight directions of OUs and thereby adjust its flight trajectory to avoid collisions.
Other UAVs (OUs). Considering the existence of airspace reuse in the low-altitude economy [36], there will also be OUs performing other tasks in the system. These UAVs may appear on the flight trajectories where the DCU collects data. To avoid collisions, the algorithm proposed in this paper considers this issue when optimizing the DCU’s flight trajectory.
Ground Equipment (GE). GE refers to devices that generate data in the low-altitude economy, such as intelligent traffic devices, power infrastructure equipment, and environmental monitoring devices [37]. These data are collected regularly by the DCU.
3.2 Models Related to UAV
This section presents models about how UAV moves, transmits data, and consumes energy in our system. The high mobility, dynamic topology, and fast-fading channels cause Doppler frequency shift [41], which induces intercarrier interference and raises the bit error rate, thereby degrading DCU data collection performance. Therefore, we use the “hover-and-fly” scheme for the DCU [32].
3.2.1 Mobility Model
The movement model is used to describe how the position of the DCU changes at times $t$ and $t+\Delta t$, that is, how the DCU moves. Given the DCU’s position ${L_{u,t}}=({x_{u,t}},{y_{u,t}},{z_{u,t}})$ and the velocity ${v_{u}}(t)=({v_{u}}{(t)_{x}},{v_{u}}{(t)_{y}},{v_{u}}{(t)_{z}})$ at time $t$, the DCU’s position at time $t+\Delta t$ is as follows:
${L_{u,t+\Delta t}}=$ (1)
$({x_{u,t}}+{v_{u}}{(t)_{x}}{t_{fly}},\ {y_{u,t}}+{v_{u}}{(t)_{y}}{t_{fly}},\ {z_{u,t}}+{v_{u}}{(t)_{z}}{t_{fly}}).$
Both DCU and GE are equipped with a single antenna [43]. The entire data collection period is divided into equal-length time slots ($\{{t_{i}}\}_{i=1}^{T/\Delta t}$). In a given time slot $\Delta t$, the process is split into two phases: a flight phase, where the DCU searches for the optimal communication position, and a data collection phase, where it hovers at that position to collect data. Therefore, given ${t_{fly}}=\Delta t/2$, ${z_{u,t}}=H$, ${v_{u}}{(t)_{z}}=0$, the distance between $L_{u,t+\Delta t}$} and $L_{u,t}$} should satisfy the following constraint:
$\left\|{L_{u,t+\Delta t}}-{L_{u,t}}\right\|\leq{v_{\max}}{t_{fly}}.$ (2)
If the DCU flies into $\mathbb{R}_{RZ}$}, the distance between $L_{u,t+\Delta t}$} and $L_{u,t}$} should satisfy the following constraint:
$\left\|{L_{u,t+\Delta t}}-{L_{u,t}}\right\|\leq{v_{{\rm{limit}}}}{t_{fly}}.$ (3)
Note that the mobility model of OUs is the same as that of DCU.
3.2.2 Data Transmission Model
The data transmission model describes how much the data collection volume of the DCU changes at times $t$ and $t+\Delta t$, that is, how much data the DCU can transmit within time $\Delta t$ ($T{D_{u,G{E_{i}}}}(t)$). Before giving how to compute $T{D_{\Delta t}}$, we first present the effective information rate of the DCU in the time slot $\Delta t$. The DCU collects data from a single randomly selected GE within its communication range in each time slot by using the standard time division multiple access [38]. When DCU collects data, it is in a hovering state without moving. The DCU begins to collect data at time $t+{t_{fly}}$ ($t_{fly}=\Delta t/2$) and lasts for ${t_{col}}=\Delta t-{t_{fly}}$. As a result, assuming the ground device that is currently collecting data is $GE_{i}$, the effective information rate during the time slot $\Delta t$ is as follows:
${B_{u,GE_{i}}}(t)=\log_{2}\left(1+SNR_{u,GE_{i}}(t)\right),$ (4)
where $SNR_{u,GE_{i}}(t)=TP_{GE_{i}}\cdot G_{u,GE_{i}}(t)/\mathcal{N}$ is the signal-to-noise ratio in the time slot $\Delta t$ beginning at time $t$, $TP_{GE_{i}}$ is the transmit power of $GE_{i}$, $G_{u,GE_{i}}(t)$ is the channel power gain from the DCU to $GE_{i}$ in the time slot $\Delta t$ beginning at time $t$, and $\mathcal{N}$ is the noise power.
Given the flight altitude of the DCU, the communication link with $GE_{i}$ is assumed to follow a line-of-sight channel model [51]. The free-space path-loss model is used to compute $G_{u,GE_{i}}(t)$ as follows [44]:
$G_{u,GE_{i}}(t)=$ (5)
[…]
where $\alpha$ is the channel power gain at $\sqrt{\left\|(x_{u,t+t_{fly}},y_{u,t+t_{fly}},H)-(x_{GE,i},y_{GE,i},0)\right\|^{2}}=1$ m distance.
Note that, if the DCU wants to transmit data, $SNR_{u,GE_{i}}(t)$ must exceed a certain threshold $\tau$ [45]. Therefore, according to Equations (4) and (5), $TD_{u,GE_{i}}(t)$ can be computed as follows:
[…]
3.2.3 Energy Consumption Model
The energy consumption model is used to describe how much the remaining energy of the the DCU changes at times $t$ and $t+\Delta t$, that is, how much energy the DCU consumes within time $\Delta t$ ($E(t)$). The energy consumption of the DCU within $\Delta t$ beginning at time $t$ can be divided into two parts: $e_{\text{fly}}(t)$ and $e_{\text{hover}}(t)$. For $e_{\text{fly}}(t)$, it can be computed as follows [47]:
$e_{\text{fly}}(t)=$ (7)
$\Big{(}\frac{\lambda}{8}\cdot\rho\eta A_{r}\Im^{3}r_{\text{rotor}}^{3}\left(1+\frac{3v_{u}(t)^{2}}{U_{\text{tip}}^{2}}\right)+\frac{1}{2}d_{0}\rho sA_{r}v_{u}(t)^{3}\Big{)}\cdot t_{\text{fly}},$
where $\lambda$ denotes the profile drag coefficient, $A_{r}$ the rotor disc area, $\Im$ the blade angular velocity, $r_{\text{rotor}}$ the rotor radius, $d_{0}$ denotes the fuselage drag coefficient, $\rho$ the density of air, $\eta$ the rotor solidity factor, $U_{\text{tip}}$ the tip velocity of the rotor blade, $\lambda$ the average induced velocity during hovering, and $v_{u}$ the velocity of DCU.
For $e_{\text{hover}}(t)$, it can be computed as follows:
$e_{\text{hover}}(t)=$ (8)
$\left((1+\Re)w^{3/2}\sqrt{1+\frac{v_{u}(t)^{4}}{4v_{0}^{4}}}-\frac{v_{u}(t)^{2}}{2v_{0}}\middle/\sqrt{2\rho A_{r}}\right)\cdot t_{\text{hover}},$
where $\Re$ is the incremental correction factor to induced power, $w$ the weight of DCU, and $v_{0}$ the mean rotor induced velocity in hover.
Note that since the energy consumed by DCU for communication is much less than that consumed for hovering [46], when calculating $e_{\text{hover}}(t)$, we ignore the energy consumed by DCU for data transmission. Besides, the UAV-related parameter values in this section follow the settings provided in [47].
4 Optimization Problem
In this section, we present the main requirements behind the optimization problem to highlight the motivation before formally defining it. Then we detail the formalized problem and its characteristic.
4.1 Requirement Analysis
The optimization problem involves two primary requirements. The first requirement includes the constraints imposed by the physical environment, and the second one involves the task-specific requirements associated with data collection in the low-altitude economy. For the former, the following requirement is taken into consideration:
R1) Limited Mobility. The moving distance of DCU during each time slot should satisfy the Equation  (2).
For the latter, the following requirements are taken into consideration:
R2) Robust Obstacle Avoidance. Since data collection in the low-altitude economy is concentrated in urban areas and there is airspace reuse in low-altitude areas [36], when the DCU conducts data collection, it not only needs to avoid ${\mathbb{R}}_{BZ}$}, ${\mathbb{R}}_{NFZ}$}, but also OUs.
R3) Compliance. Unlike general data collection tasks, DCU operating in urban low-altitude environments must comply with strict regulations. They must reduce speed near to minimize noise and meet regulatory standards.
R4) Economy. Within the low-altitude economy, data collection tasks must jointly consider the maximization of data acquisition and the minimization of energy consumption to maximize the data collection task benefits. Consequently, the DCU is required not only to collect sufficient data but also to complete the task along an energy-efficient trajectory that minimizes energy costs.
4.2 Problem Formulation
According to the requirement analysis introduced in Section 4.1, we try to optimize the flight trajectory of DCU while maximizing the amount of collected data. The optimization problem can be formally formulated as follows:
$\textbf{OP:}\quad\arg\max\sum\nolimits_{t=0}^{T}TD_{u,GE_{i}}(t),$ (9)
s.t.
$GE_{i}=\arg\max\left(\{B_{u,GE_{i}}(t)\}_{i=1}^{N_{GE}}\wedge SNR_{u,GE_{i}}(t)\geq\tau\right),$ (9a)
$t+1=t+\Delta t,\quad\forall t,$ (9b)
$L_{u,t+1}=(x_{u,t}+v_{u}(t)_{x}\cdot t_{\text{fly}},\ y_{u,t}+v_{u}(t)_{y}\cdot t_{\text{fly}},\ z_{u,t}),$ (9c)
$v_{u}(t)\leq v_{\max},\quad\forall t,$ (9d)
$v_{u}(t)\leq v_{\text{limit}},\quad\text{if }L_{u,t}\in\{\mathbb{R}_{RZ,m}\}_{m=1}^{N_{RZ}},$ (9e)
$\begin{aligned} &(L_{u,t},L_{u^{\prime},t})\notin\{\mathbb{R}_{BZ,n}\}_{n=1}^{N_{BZ}}\\ &\wedge\ (L_{u,t},L_{u^{\prime},t})\notin\{\mathbb{R}_{NFZ,l}\}_{l=1}^{N_{NFZ}},\quad\forall t,\end{aligned}$ (9g)
$\left\|L_{u,t}-L_{u^{\prime}_{j},t}\right\|>0,\quad\forall t,\ j\in[1,N_{U^{\prime}}],$ (9h)
$\sum\nolimits_{t=0}^{T}(e_{\text{fly}}(t)+e_{\text{hover}}(t))\leq E_{\text{limit}}.$ (9i)
The objective function OP maximizes the amount of data collected over the given time slots. Constraint (9a) selects a GE to collect data because the DCU follows a maximum signal strength strategy to connect to the GE with the highest received signal power [51], while (9b) presents the relationship between $t$ and $t+1$ according to the data transmission model. Constraint (9c) governs the position updates of the DCU according to the mobility model. Constraint (9d) enforces the maximum speed limit, and Constraint (9e) enforces the maximum speed limit in $\mathbb{R}_{RZ}$ to satisfy R1 and R3. Constraints (9f), (9g), and (9h) ensure that the DCU operates within $\mathbb{A}$, avoids $\mathbb{R}_{NFZ}$, and prevents collisions with $\mathbb{R}_{BZ}$ or OUs to satisfy R2 and R3. Constraint (9i) ensures the maximum energy consumption of the DCU in a data collection task to satisfy R4.
4.3 Problem Characteristics
This optimization problem involves continuous control, nonlinear constraints, dynamic environmental factors (such as OUs), as well as regulatory and energy limitations, making it structurally complex and difficult. Traditional optimization algorithms often perform inefficiently and lack adaptability when dealing with such high-dimensional, non-convex, dynamic, and partially observable problems. In contrast, DRL offers adaptive learning capabilities, making it well-suited for continuous control and long-term optimization. LLMs, on the other hand, can provide rule-compliant and reasoning decisions in critical scenarios by leveraging rich prior knowledge. The integration of DRL and LLM leverages efficiency, generalization, and safety, making it more suitable than traditional methods to solve complex UAV trajectory planning problems.
5 Proposed Algorithm
This section first formulates the optimization problem from Section 4 as a POMDP to enable solution via DRL. We then present SAC algorithm, which serves as the foundation for our algorithm. Finally, we introduce the proposed algorithm and demonstrate its effectiveness in addressing the formulated problem.
5.1 POMDP Modeling
As the future position of the DCU is determined solely by its current state (i.e., position and velocity), the trajectory planning problem satisfies the Markov property. Nevertheless, due to the DCU’s limited ability to perceive the full environmental state, the problem is more appropriately formulated as a POMDP, defined by a corresponding tuple $(\mathcal{O},\mathcal{A},r,\mathcal{P},\delta)$, where the observation space $\mathcal{O}$ replaces the state space [48]. Then, the details of $\mathcal{O}$, $\mathcal{A}$, and $r$ in $(\mathcal{O},\mathcal{A},r,\mathcal{P},\delta)$ are as follows:
Observation Space $\mathcal{O}$: $o_{t}\in\mathcal{O}$ is the state that the DCU can observe at time $t$, where $o_{t}=(o_{GE,t},o_{Z,t},o_{u^{\prime},t},o_{u,t}).$
  • $o_{GE,t}$ is the state about all GEs and $o_{GE,t}=\{L_{GE,i},DV_{GE_{i},t},TP_{GE,i}\}_{i=1}^{N_{GE}}$.
  • $o_{Z,t}$ is the state about all zones and $o_{Z,t}=\{\{\mathbb{R}_{NFZ,l}\}_{l=1}^{N_{NFZ}},\{\mathbb{R}_{RZ,m}\}_{m=1}^{N_{RZ}},\{\mathbb{R}_{BZ,n}\}_{n=1}^{N_{BZ}}\}$.
  • $o_{u^{\prime},t}$ is the state about all OUs and $o_{u^{\prime},t}=\{L_{u^{\prime}_{j},t},v_{u^{\prime}_{j},t}\}_{j=1}^{N_{U^{\prime}}}$.
  • $o_{u,t}$ is the state about DCU itself and $o_{u,t}=\{L_{u,t},v_{u,t},E(t),TD_{u,GE_{i}}(t),\mathbb{T}\mathbb{A},\mathbb{L}\mathbb{A}\}$.
Action Space $\mathcal{A}$: $a_{t}\in\mathcal{A}$ is the sampled action at time $t$ and $a_{t}=(v_{x}(t),v_{y}(t),0),$ which means the DCU will move at the velocity $(v_{x}(t),v_{y}(t),0)$ at time $t$.
Reward function $r$: $r$ is used to effectively guide the trajectory planning during the data collection process and plays a critical role in both the efficiency and performance of policy learning. To facilitate learning an optimal policy aligned with the objective function OP and its associated constraints, we formulate a composite reward function, expressed as follows:
$r_{t}=r_{1t}+r_{2t}+r_{3t}+r_{4t}+r_{5t}+r_{6t}+r_{7t}+r_{8t}+r_{9t}.$ (10)
$r_{1t}$} encourages the DCU to collect data, and is as follows:
${r_{1t}}=\left\{\begin{array}[]{ll}{\sigma_{1}},&\text{if }TD_{u,GE_{i}}(t+1)\geq 0,\\ 0,&\text{otherwise}.\end{array}\right$.
$r_{2t}$} is used to ensure that the DCU will not collide with OUs and is as follows:
${r_{2t}}=\left\{\begin{array}[]{ll}0,&\text{if }d_{t,u^{\prime}}\geq d_{safe,u^{\prime}},\\ -\sigma_{2}\times\frac{d_{safe,u^{\prime}}-d_{t,u^{\prime}}}{d_{safe,u^{\prime}}-d_{\min,u^{\prime}}},&\text{if }d_{\min,u^{\prime}}<d_{t}<d_{safe,u^{\prime}},\\ -\sigma_{2},&\text{if }d_{t,u^{\prime}}\leq d_{\min,u^{\prime}},\end{array}\right$.
where ${d_{t,u^{\prime}}}=\arg\min\{||L_{u,t}-L_{u^{\prime}_{j},t}||\}_{j=1}^{N_{U^{\prime}}}$, $d_{\min,u^{\prime}}$} is the minimum safe distance between the DCU and OUs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,u^{\prime}}$} is the safe distance threshold between the DCU and OUs.
$r_{3t}$} is used to ensure that the DCU will not collide with tall buildings and is as follows:
[…]
where ${d_{t,BZ}}=\arg\min\{||L_{u,t}-\mathbb{R}_{BZ,n}||\}_{n=1}^{N_{BZ}}$, $d_{\min,BZ}$} is the minimum safe distance between the DCU and tall buildings, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,BZ}$} is the safe distance threshold between the DCU and tall buildings.
$r_{4t}$} is used to ensure that the DCU will not fly into NFZs and is as follows:
$r_{4t}=$ (14)
[…]
where $d_{t,NFZ}=\arg\min\left\{\|L_{u,t}-\mathbb{R}_{NFZ,l}\|\right\}_{l=1}^{N_{NFZ}}$, $d_{\min,NFZ}$ is the minimum safe distance between the DCU and NFZs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,NFZ}$ is the safe distance threshold between the DCU and NFZs.
$r_{5t}$} is used to ensure that the speed of the DCU when flying into residential areas meets the speed limit and is as follows:
[…]
$r_{6t}$} is used to ensure that the DCU should not fly out of the given area and is as follows:
[…]
where $d_{t,\mathbb{A}}=\|L_{u,t}-\mathbb{A}\|$ is the distance between the DCU and the boundary of $\mathbb{A}$ at time $t$, $d_{\min,\mathbb{A}}$ is the minimum safe distance between the DCU and the boundary of $\mathbb{A}$, which is used for the DCU to adjust the speed to avoid flying out, and $d_{safe,\mathbb{A}}$ is the safe distance threshold between the DCU and the boundary of $\mathbb{A}$.
$r_{7t}$ is used to encourage that the DCU can land within the landing area and is as follows:
[…]
$r_{8t}$ is used to encourage the DCU to return to the landing area before its energy is exhausted and is as follows:
[…]
where $e_{\min}(t)=e_{\text{fly}}(t)\cdot d_{t,\mathbb{LA}}/v_{\max}$.
$r_{9t}$ is used to encourage the DCU can complete the data collection task quickly to reduce energy consumption and is as follows:
$r_{9t}=-\sigma_{9}.$ (19)
5.2 SAC Algorithm
SAC is a state-of-the-art off-policy DRL algorithm based on the maximum entropy framework [39]. Its off-policy nature allows flexible integration of external guidance from the LLM without compromising training stability or policy optimization efficiency. Additionally, SAC is well-suited for our problem, which involves a continuous action space. Although we adopt SAC in this study, we believe that other off-policy DRL algorithms supporting continuous action spaces may also achieve comparable performance, which is detailed in Section 6.
The maximum entropy DRL incorporates the policy’s entropy into the objective function [40], modifying the learning goal as follows:
$\pi^{*}=\arg\max_{\pi}\sum_{t=0}^{\infty}\mathbb{E}_{\mathcal{P}\pi}\left[\delta^{t}\left(r_{t}+\mu P(\pi(\cdot\mid o_{t}))\right)\right],$ (20)
where $\delta$ is the discounted factor and $\mu$ is the temperature parameter. The policy entropy $P(x)$ is defined as follows:
$P(x)=-\sum\nolimits_{x}p(x)\log p(x).$ (21)
Given the soft value-state function and the soft value function as follows:
$Q_{\text{soft}}(o_{t},a_{t})=r_{t}+\mathbb{E}_{\mathcal{P}\pi}\left[\sum\nolimits_{i=1}^{\infty}\delta^{i}\left(r_{t+i}+\mu P(\cdot\mid\pi)\right)\right],$ (22)
$V_{\text{soft}}(o_{t})=\mu\log\int_{\mathcal{A}}\exp\left(\frac{Q_{\text{soft}}(o_{t},a_{t+1})}{\mu}\right)da_{t+1}.$ (23)
They can be iterated in the following ways:
$Q_{\text{soft}}(o_{t},a_{t})=r_{t}+\delta\mathbb{E}_{\mathcal{P}}\left[V_{\text{soft}}(o_{t+1})\right],$ (24)
$V_{\text{soft}}(o_{t})=\mathbb{E}_{\pi}\left[Q_{\text{soft}}(o_{t},a_{t})-\mu\log\pi(a_{t}\mid o_{t})\right].$ (25)
Therefore, the policy $\pi$ is as follows:
$\pi(o_{t},a_{t})=\exp\left(\frac{Q_{\text{soft}}(o_{t},a_{t})}{\delta}-V_{\text{soft}}(o_{t})\right).$ (26)
In SAC, both $\pi$ and $Q_{\text{soft}}$ are parameterized using neural networks. This enables SAC to handle continuous state and action spaces, making it well-suited for solving complex problems and enhancing controller performance in scenarios, such as trajectory planning. The neural networks associated with $Q_{\text{soft}}$ are updated at each step by minimizing the following loss function:
$J_{Q}(\varphi_{i})=\mathbb{E}_{(o_{t},a_{t},r_{t},o_{t+1})\sim D}\left[\left(Q_{i}^{\theta}(o_{t},a_{t})-y(r_{t},o_{t+1})\right)^{2}\right],$ (27)
where $y(r_{t},o_{t+1})$ is as follows:
$y(r_{t},o_{t+1})=$ (28)
$r_{t}+\delta\left(\min_{j=1,2}Q_{i}^{\theta_{\text{tar}}}(o_{t+1},\tilde{a}_{t+1})-\mu\log\pi^{\phi}(\tilde{a}_{t+1}\mid o_{t+1})\right).$
For $\pi$, with the sampled action, it is optimized as follows:
$J_{\pi}(\phi)=$ $\mathbb{E}_{o_{t}\sim D,\varsigma\sim N}\big{[}\mu\log\pi^{\phi}(\tilde{a}_{t}^{\phi}(o_{t},\varsigma)\mid o_{t})$- (29)
$\min_{j=1,2}Q_{i}^{\theta_{\text{tar}}}(o_{t},\tilde{a}_{t}^{\phi}(o_{t},\varsigma))\big{]},$
where $\tilde{a}_{t}^{\phi}(o_{t},\varsigma)$ is as follows:
$\tilde{a}_{t}^{\phi}(o_{t},\varsigma)=\tanh\left(\beta^{\phi}(o_{t})+\alpha^{\phi}(o_{t})\odot\varsigma\right),\quad\varsigma\sim\mathcal{N}(0,1),$ (30)
where $\beta$ and $\alpha$ represent the mean and standard deviation of the Gaussian distribution, respectively; $\varsigma$ denotes the noise; and $\mathcal{N}(0,1)$ indicates the standard normal distribution.
Figure 2: The proposed algorithm integrates an LLM into SAC for conditional action selection. The LLM is invoked to generate actions when obstacles are detected; otherwise, actions are selected by the SAC policy network. The chosen actions interact with the environment, and transitions are stored to train the critic networks.
5.3 Our Algorithm
This section presents the core design principles of our trajectory planning algorithm, outlines its detailed implementation, and provides a comprehensive analysis of its computational complexity.
5.3.1 Main Idea
Based on the POMDP formulation in Section 5.1 and the continuous nature of DCU’s velocity, we adopt SAC algorithm to optimize the DCU’s flight trajectory during the data collection task. However, in the context of low-altitude economy scenarios, where strict regulatory compliance and robust obstacle-avoidance capability are critical, the use of DRL alone may be insufficient due to its reliance on trial-and-error learning and lack of prior domain knowledge. To address this issue, we propose an enhanced decision-making algorithm by integrating an LLM into the action selection module, as illustrated in Fig.  2. Specifically, when the DCU detects the presence of nearby obstacles (e.g., BZs, NFZs, RZs, or OUs), a decision condition is triggered that activates the LLM-based action selection pathway. In this mode, a Prompt Generator compiles semantic information, including environmental context, DCU status, and obstacle details, into a structured prompt, which is then input to an LLM. The LLM responds with a recommended action, leveraging its rich prior knowledge to ensure safety and regulatory compliance. An Extractor then parses the LLM’s response into actionable control parameters for the DCU. Conversely, when no immediate obstacle is detected, the system defaults to the standard SAC policy network for action selection. This hybrid structure enables the agent to benefit from both data-driven learning and knowledge-driven reasoning: the SAC component provides efficient trajectory planning in general scenarios, while the LLM supplements decision-making in critical or unfamiliar environments requiring sophisticated judgment. By combining the generalization capability of DRL with the reasoning capability and prior knowledge of an LLM, the proposed algorithm achieves superior performance in complex real-world scenarios, ensuring not only energy efficiency but also operational safety and compliance.
5.3.2 Algorithm Detail
Our proposed algorithm, which is detailed in Algorithm 1, is designed to enable the DCU to effectively perform data collection by learning an optimal velocity control policy that guides its trajectory to maximize data acquisition and avoid collision. The main process of our algorithm is shown in Fig. 2.
The proposed algorithm takes the environment as input (Line 1) and outputs an optimal policy network (Line 2) that governs the DCU velocity control. Upon initialization (Line 3), the initialization of the Q-function parameters, policy network parameters, the LLM, the experience pool, and the target network parameters is conducted.
For each training episode, the environment is first reset (Line 5). Then, at each time step within the episode, the algorithm observes the current state and determines whether to invoke the LLM for action selection (Lines 7–8). It is important to note that this state ($\tilde{o}_{t}$) differs from the state ($o_{t}$) used by the policy network. Specifically, it refers to the pre-processed, semantically rich environmental information, such as the positions of obstacles within the DCU’s observation range, as illustrated in Fig. 2. If the decision condition is met, i.e., there are obstacles near the DCU, the LLM is utilized to select an action for the DCU based on the semantic state information generated by the prompt generator. Otherwise, the action is selected using the policy network (Lines 9–12). The selected action is then executed, resulting in the next state and corresponding reward for the subsequent time slot. Simultaneously, the current state, action, and reward are stored in the experience pool (Lines 14–15). The algorithm then checks whether the current episode has terminated. If so, it proceeds to the next episode (Line 17).
For each gradient update step, a batch of samples is drawn from the experience pool (Line 21). Based on this batch, the target values are computed, and the Q-function, policy network, and target network parameters are updated accordingly (Lines 22–24). This process repeats until the episode concludes, after which the next episode begins. Upon completion of all training episodes, the algorithm outputs the final optimized policy network.
Algorithm 1 SAC with LLM-based Trajectory Planning Algorithm
1:Input: The environment $\mathcal{E}$
2:Output: The optimal policy network $\pi^{*}$
3:Initialize: Q-function parameters $\theta_{1},\theta_{2}$, policy parameter $\phi$, the LLM $\pi^{LLM}$, experience pool $D$, target parameters $\theta_{1,\text{tar}},\theta_{2,\text{tar}}$
4:for $\text{episode}=1,2,\ldots,L$ do
5:  Initialize environment: $\mathcal{E}\leftarrow\text{EnvReset()}$
6:  for each $\text{step}=1,2,\ldots,N$ do
7:   Observe state: $o_{t}\leftarrow\text{ObsState}(\mathcal{E})$
8:   Judge if using the LLM: $\text{bool}\leftarrow\text{useLLM}(\tilde{o}_{t})$
9:   if bool then
10:     Select action from the LLM: $a_{t}\leftarrow\pi^{LLM}(\tilde{o}_{t})$
11:   else
12:     Select action from policy network: $a_{t}\leftarrow\pi^{\phi}(o_{t})$
13:   end if
14:   Execute action and get new state and reward: $(o_{t+1},r_{t},\text{done})\leftarrow\text{ExeAct}(a_{t})$
15:   Store transition in experience pool: $(o_{t},r_{t},a_{t},o_{t+1})\to D$
16:   if done then
17:     Reset the environment and continue
18:   end if
19:  end for
20:  for each gradient step do
21:   Randomly sample a batch: $(o_{t},r_{t},a_{t},o_{t+1})\leftarrow D$
22:   Compute targets and update Q-functions: $Q_{1}^{\theta},Q_{2}^{\theta}$
23:   Update policy network $\pi^{\phi}$
24:   Update target parameters: $\theta_{1,\text{tar}},\theta_{2,\text{tar}}$
25:  end for
26:end for
5.3.3 Algorithm Computational Complexity Analysis
During training, our algorithm largely follows the SAC framework for policy optimization. However, unlike standard SAC, our algorithm intermittently replaces the policy network’s action outputs with those generated by an external LLM based on specific conditions. As a result, the overall training complexity consists of two components: the standard SAC update steps and the LLM inference overhead incurred during action selection. Due to each SAC training step including policy network updates with complexity $O(d_{\pi})$, Q-network updates with complexity $O(d_{Q})$, and target network updates with negligible computational cost, the training computation complexity of SAC is $O(n\cdot(d_{\pi}+2d_{Q})),$ where $n$ is the batch size [52]. Assuming the LLM is queried for $p\cdot n$ samples per batch, the training computation complexity of the LLM is $O(p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}),$ where $L_{\text{LLM}}$ is the number of tokens and $d_{\text{LLM}}$ is the LLM hidden layer size [54]. Therefore, the training computation complexity of our algorithm is $O((1-p)\cdot n\cdot(d_{\pi}+2d_{Q})+p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}).$ Despite the additional overhead introduced by occasional LLM queries, the overall computation complexity remains comparable to standard SAC, and the selective use of LLM ensures that the efficiency of the training process is largely preserved, offering a favorable trade-off between performance and cost.
6 Experiments
In this section, we present the experimental results of the proposed algorithm. Prior to that, we describe the experimental setup and evaluation metrics.
6.1 Experiment Setting
For the hardware, all experiments in this study were conducted on a computing platform equipped with Intel® Xeon® Gold 6230 CPUs and NVIDIA GeForce RTX 4090 GPUs.
TABLE III: Environment Parameters Setting
Para. Value Para. Value Para. Value
$N_{GE}$ [11, 15] $v_{\max}$ 10 m/s $\mathcal{N}$ $10^{-6}$
$DV_{GE_{i}}$ [1, 3] $v_{\text{limit}}$ 5 m/s PR 20 m
$TP_{GE,i}$ 0.01 W $E_{\text{total}}$ $1\times 10^{6}$ J $d_{\min},d_{\text{safe}}$ 5 m, 15 m
$N_{U^{\prime}}$ [3, 7] $E_{\text{limit}}$ $8\times 10^{5}$ J $\tau$ 3.6
6.1.1 Environment Parameters
The experimental environment is simplified to a two-dimensional space with $\mathbb{A}$ of 500 m in both length and width. $\mathbb{T}\mathbb{A}$ and $\mathbb{L}\mathbb{A}$ are square regions measuring $25\times 25$ m, located at the top-left and bottom-right corners of $\mathbb{A}$, respectively. Given the urban deployment scenario, the environment includes rectangular ${\mathbb{R}}_{NFZ}$}, ${\mathbb{R}}_{RZ}$}, and ${\mathbb{R}}_{BZ}$} with the range of [50, 200] m for both length and width. The number of such regions are randomly chosen from the range [1, 3]. The weights assigned to the components of the reward function are set to [8, 0.4, 0.4, 0.4, 0.6, 0.4, 0.6, 1, 0.01][8,\ 0.4,\ 0.4,\ 0.4,\ 0.6,\ 0.4,\ 0.6,\ 1,\ 0.01] and can be flexibly adjusted to reflect different task priorities. Other environmental parameters are detailed in Table III.
6.1.2 Algorithm Parameters
For the algorithm, hyperparameters used by the proposed algorithm and other baselines, including SAC+LLM, SAC, Proximal Policy Optimization (PPO) [55], and DDPG [56], are shown in Table IV. For the LLM used in our algorithm and other baselines, including LLM and SAC+LLM, we choose the DeepSeek-R1, which has a total of 672B parameters and 37B activated parameters [49].
TABLE IV: Hyperparameters Setting
Parameter Value
Learning rate of actor network 0.0001
Learning rate of critic network 0.0003
Number of episodes 4000
Clip range 0.2
Batch size 256
Discount factor 0.99
Number of hidden layers of DNNs 2
Hidden layer size 64
6.2 Evaluation Metrics
For the metrics used to evaluate the proposed algorithm and other baselines, we use the metrics commonly adopted in general UAV data collection trajectory planning tasks [50]. In addition, we include metrics that reflect the specific requirements in this paper. All metrics are directly aligned with the core demands of urban low-altitude data collection in the emerging low-altitude economy, where robust obstacle avoidance (R2), compliance (R3), and economy (R4), mentioned in Section 4.1, are essential. Assuming $N_{\text{total}}$ is the number of experimental tests:
  • Data collection rate (DCR): This metric quantifies the average ratio of successfully collected data by the DCU relative to the total target data volume in $N_{\text{total}}$ data collection tasks:
    $\text{DCR}=\frac{\sum\nolimits_{i=1}^{N_{\text{test}}}D_{\text{DCU},i}}{\sum\nolimits_{i=1}^{N_{\text{test}}}D_{\text{total},i}},$ (35)
    where $D_{\text{DCU},i}$ is the collected data volume and $D_{\text{total},i}$ is the total target data volume in the $i$-th task.
  • Collision rate (CR): This metric quantifies the ratio of data collection tasks in which the DCU collides with OUs, $\mathbb{R}_{BZ}$, or flies into $\mathbb{R}_{NFZ}$:
    $\text{CR}=\frac{N_{C}}{N_{\text{total}}},$ (36)
    where $N_{C}$ is the number of tasks involving collisions or regulatory violations.
  • Successful landing rate (SLR): This metric quantifies the ratio of data collection tasks in which the DCU can return to $\mathbb{LA}$ before depleting its energy:
    $\text{SLR}=\frac{N_{\text{SLR}}}{N_{\text{total}}},$ (37)
    where $N_{\text{SLR}}$ is the number of tasks where the DCU successfully lands.
In addition to the commonly used metrics mentioned above, the following indicators are also employed to evaluate the effectiveness of satisfying the requirement for data collection trajectory planning under the low-altitude economy.
  • Regulatory violation rate (RVR): This metric quantifies the ratio of tasks in which the DCU exceeds speed limits within $\mathbb{R}_{RZ}$:
    $\text{RVR}=\frac{N_{\text{RVR}}}{N_{\text{total}}},$ (38)
    where $N_{\text{RVR}}$ is the number of such violations.
  • Energy consumption rate (ECR): This metric quantifies the average ratio of energy consumed by the DCU relative to the total energy:
    $\text{ECR}=\frac{\sum\nolimits_{i=1}^{N_{\text{test}}}E_{\text{ECR},i}}{\sum\nolimits_{i=1}^{N_{\text{test}}}E_{\text{total},i}},$ (39)
    where $E_{\text{ECR},i}$ is the consumed energy and $E_{\text{total},i}$ is the total available energy in the $i$-th task.
6.3 Experiment Results
In this section, we present a comparative analysis between the proposed algorithm  [^1] The source code is available at: https://github.com/BJTU-STIC/DRL-LLM and several baseline methods in terms of the DCU flight trajectories and the evaluation metrics mentioned above. The parameter settings of these baselines are introduced in Section 6.1.2. The baselines are as follows:
  • Heuristic method: A rule-based strategy is used to determine the actions of the DCU for data acquisition and obstacle avoidance. The DCU sequentially collects data from the nearest GE until all data is acquired. If a static obstacle such as $\mathbb{R}_{NFZ}$ or $\mathbb{R}_{BZ}$ is encountered, the DCU performs detouring maneuvers. For dynamic obstacles like OUs, it either hovers in place or retreats to avoid collision until the OU moves away. To limit the noise impact when flying over $\mathbb{R}_{RZ}$, the DCU must reduce its speed below $v_{\text{limit}}$ in accordance with regulatory requirements [6]. Furthermore, if the remaining energy is only sufficient for returning to $\mathbb{LA}$, the DCU will immediately head back.
  • LLM: An LLM, DeepSeek-R1, is used in real-time to decide the actions of the DCU. We provide the LLM with the state information within a 50 m radius around the DCU, as well as the positions of the two nearest GEs. In addition, we also provide the LLM with $E_{\text{limit}}$ of the DCU.
  • SAC+LLM (Fixed-$fp$): SAC is used during training, with a certain probability ($fp\in\{0.3,0.4,0.5,1.0\}$) of delegating action selection to the LLM instead of the policy network. The LLM used here is the same as the one employed in the proposed algorithm, i.e., DeepSeek-R1.
  • SAC: A pure deep reinforcement learning approach using the SAC without LLM integration.
  • PPO: A representative on-policy reinforcement learning method.
  • DDPG: A classic off-policy actor-critic method designed for continuous control.
Figure 3: Training reward curves for the proposed algorithm and baseline algorithms over 4000 episodes, including SAC, PPO, DDPG, and SAC+LLM variants with different invocation probabilities.
(a) (a1)
(b) (a2)
(c) (a3)
(d) (a4)
(e) (b1)
(f) (b2)
(g) (b3)
(h) (b4)
(i) (c1)
(j) (c2)
(k) (c3)
(l) (c4)
(m) (d1)
(n) (d2)
(o) (d3)
(p) (d4)
(q) (e1)
(r) (e2)
(s) (e3)
(t) (e4)
Figure 4: Representative snapshots of DCU flight trajectories during data collection under different algorithms. (a1-a4): Heuristic; (b1-b4): LLM; (c1-c4): SAC+LLM (0.3); (d1-d4): SAC; (e1-e4): Ours. The trajectories illustrate obstacle avoidance, data collection behavior, and regulation compliance. PPO, DDPG, and SAC+LLM (1.0) are excluded for clarity, as their policies either fail to converge or produce unstable behaviors with limited comparative value. Obstacles (NFZs, BZs, RZs, and OUs), and GEs are highlighted in each subplot to illustrate the algorithms’ compliance and avoidance behavior.
6.3.1 Flight Trajectories about Baselines and Our Algorithm
Before presenting the flight trajectories, we first provide the training reward curves for the baselines related to DRL and the proposed algorithm. As illustrated in Fig. 3, these curves reflect the convergence behavior of each algorithm over 4000 training episodes. The proposed algorithm achieves the highest and most stable cumulative rewards, demonstrating superior policy quality and learning efficiency. Both SAC and PPO exhibit clear convergence trends, with PPO showing slightly slower but stable improvement. DDPG also improves steadily but converges to a lower reward plateau, suggesting limited effectiveness in handling the complex constraints of this environment. In contrast, the SAC+LLM variants with randomly invoked LLMs (at probabilities 0.3, 0.4, 0.5, and 1.0) exhibit significantly lower and more unstable reward trajectories. Notably, the SAC with random LLM 1.0 variant, where the LLM selects all actions, fails to converge effectively, highlighting the detrimental impact of fully replacing policy learning with LLM reasoning during training. This further supports our core design: selectively invoking the LLM in safety-critical contexts ensures that high-level reasoning complements rather than disrupts the DRL policy learning process. Overall, these results validate that our hybrid algorithm balances learning efficiency and decision safety, outperforming both standalone DRL methods and indiscriminate LLM integration strategies.
Fig. 4 illustrates segments of the DCU’s flight trajectories during data collection under both the proposed algorithm and the baselines. We highlight four typical stages: static obstacle avoidance, data collection paths, dynamic obstacle avoidance, and return to $\mathbb{LA}$. While PPO, DDPG, and SAC+LLM (0.4, 0.5, 1.0) are included in the training reward comparison (Fig. 3) to evaluate learning performance, their flight trajectories are not shown in the subsequent visualizations. This is because these methods either fail to converge to stable policies or yield suboptimal control behaviors that do not contribute additional insights beyond those already observed from SAC and SAC+LLM (0.3) variants. To maintain clarity and focus, we present representative trajectory comparisons only for the most informative and competitive baselines. Figs. 4(a1) and (b1) illustrate that both the heuristic and LLM approaches predominantly follow a “nearest GE” strategy for data collection. However, the LLM approach exhibits more efficient trajectory planning, demonstrating a tendency to select shorter and more obstacle-aware trajectories. Notably, Fig. 4(b2) indicates that the LLM does not rigidly adhere to the nearest GE principle; instead, it appears to balance the trade-off between obstacle avoidance and data acquisition efficiency. Despite this, as shown in Figs. 4(a4) and (b4), adherence to the nearest GE strategy tends to increase the overall flight distance. As a result, both the heuristic and the LLM approaches often encounter a scenario where the DCU reaches its energy limit before completing the data collection task, thereby necessitating an early return to $\mathbb{LA}$. Fig. 4(c) presents the performance of SAC+LLM, wherein the LLM is stochastically invoked during training to determine the DCU’s actions. This integration does not effectively enhance SAC’s decision-making capability and, in fact, leads to performance degradation. As observed in Figs. 4(c1) and (c4), this method results in suboptimal obstacle avoidance and frequently prevents the DCU from successfully returning to $\mathbb{LA}$, often exhausting its energy mid-flight. Figs. 4(d) and (e) depict the flight trajectories generated by the SAC algorithm and the proposed algorithm, respectively. From Figs. 4(d1) and (e1), it can be observed that SAC adopts a more conservative obstacle avoidance strategy, which inadvertently increases flight distance. Figs. 4(d2) and (e2) show that SAC may overlook distant GEs, resulting in reduced data acquisition. Furthermore, Figs. 4(d3) and (e3) indicate that SAC tends to omit data collection opportunities in the presence of obstacles, adversely affecting overall data completeness. Finally, Figs. 4(d4) and (e4) reveal that SAC may opt to traverse residential areas directly. Empirical testing suggests that SAC occasionally exceeds speed limits within these zones, thereby compromising compliance with operational regulations.
6.3.2 Metrics about Baselines and Our Algorithm
The experimental results in Figs. 5–10 comprehensively assess five critical performance metrics, DCR, CR, SLR, RVR, and ECR, under varying numbers of OUs, GEs, NFZs, BZs, and RZs. We analyzed various algorithms for the requirements: robust obstacle avoidance (R2), compliance (R3), and economy (R4). This further demonstrates how the algorithm in this paper addresses Challenges 1–3 mentioned in Section 1.
Fig. 5 presents the performance of different algorithms under increasing numbers of OUs, representing dynamic obstacles. Our algorithm consistently outperforms all baselines, maintaining the highest DCR and the lowest CR across all conditions, while also achieving the most energy-efficient trajectories (lowest ECR). Notably, while SAC performs reasonably well, it suffers from rising CR as $N_{U^{\prime}}$ increases. SAC+LLM achieves high DCR but exhibits extremely poor robustness, with CR exceeding 93%, indicating that the naive LLM integration without appropriate control undermines safety. DDPG and PPO offer moderate DCR improvements over heuristic and LLM methods, but their CRs remain higher and ECRs less efficient than ours. Notably, the standalone LLM method maintains a low CR at first, but both its DCR and ECR remain inferior to learning-based approaches, and CR increases sharply under denser conditions. These results highlight our algorithm’s superiority in dynamic, uncertain environments (R2, Challenges 1&3), and its capability to maintain economic efficiency (R4, Challenge 2) without compromising safety.
Fig. 6 evaluates the algorithms under increasing numbers of GEs, reflecting higher task complexity. Our algorithm consistently achieves the highest DCR (DCR $\approx$ 1.0 and near-perfect SLR (SLR $\approx$ 1.0), demonstrating mission reliability. While SAC also performs well, its DCR and SLR decline slightly with complexity. DDPG and PPO achieve moderate DCR and SLR across all GE levels, outperforming heuristic and LLM methods in data collection yet trailing behind SAC and our algorithm in both reliability and safety. Their CRs remain relatively higher than our algorithm and SAC, indicating less robustness when handling complex tasks. SAC+LLM exhibits high DCR at low complexity but suffers a sharp drop in SLR (to as low as 5.63%) and persistently high collision risk, indicating poor robustness in full-mission execution. The LLM alone maintains stable SLR, but its DCR degrades significantly as targets increase, limiting its effectiveness. These results highlight our algorithm’s ability to plan efficient and reliable trajectories in increasingly complex environments, addressing Challenge 2 and Challenge 3.
Figs. 7 and  8 examine algorithmic performance under increasing numbers of NFZs and BZs, both representing static obstacles. Our algorithm maintains DCR $\approx$ 1.0 and minimal CR (CR $\approx$ 0), even as obstacle density rises, clearly demonstrating superior robust static obstacle avoidance (R2). In contrast, SAC+LLM suffers from a sharp increase in CR despite relatively high DCR, indicating that the LLM component alone is insufficient to compensate for suboptimal DRL decisions in constrained spaces. The LLM alone maintains zero CR at low obstacle counts but with consistently low DCR, reflecting limited planning capability. SAC shows moderate DCR but increasing CR (up to 15.67%), suggesting vulnerability to static obstacles. DDPG and PPO demonstrate stronger performance than SAC+LLM and LLM, achieving relatively high DCR. However, both suffer from noticeably higher CR compared to SAC and our algorithm, especially as obstacle density increases, which reflects limited robustness in handling constrained static environments. Their ECRs are moderate, indicating some economic trajectory planning but still lagging behind our algorithm. Furthermore, our algorithm consistently achieves the lowest ECR across all conditions, reinforcing its ability to optimize trajectories economically. These results affirm the effectiveness of our algorithm in addressing Challenge 1 and partially Challenge 2 (maintaining the economy).
Fig. 9 explores the algorithms’ ability to balance task completion and compliance under increasing numbers of RZs, regions where entry is permissible but potentially non-compliant. Our algorithm consistently achieves the highest DCR while maintaining low CR and RVR, showcasing strong capability in resolving Challenge 3. SAC+LLM, although maintaining high DCR, incurs both high CR and RVR, suggesting its reliance on the LLM is insufficient to enforce compliance under ambiguity. SAC alone shows moderate RVR but suffers from increasing CR. DDPG and PPO also demonstrate decent task completion with relatively high DCR. However, both algorithms exhibit noticeably higher CR and RVR compared to SAC and our algorithm, indicating their limited ability to handle ambiguous regulatory boundaries. While they outperform SAC+LLM in compliance, their safety and adherence still fall short in soft-constrained environments. In contrast, the LLM and heuristic methods avoid almost all violations (0.34% and 0%), but their performance stagnates at low DCR (74.62% and 76.92%), implying over-conservatism and suboptimal data acquisition. Overall, our algorithm uniquely achieves both mission effectiveness and regulatory adherence, excelling in environments with soft constraints.
The radar chart in Fig. 10 clearly demonstrates the superior performance of our proposed algorithm across all five metrics. It achieves high DCR and low ECR, indicating effective data collection and energy-efficient mission completion (R4). Simultaneously, it maintains high SLR and low CR and RVR, ensuring both obstacle avoidance and regulatory compliance, critical for safe operation in urban, airspace-reused environments (R2 and R3).
In summary, the results validate that our algorithm, by selectively engaging the LLM for obstacle avoidance, empowers the DCU to make informed and adaptive decisions in complex, dynamic environments. This targeted integration enables our algorithm to uniquely balance the three core demands of the low-altitude economy: R2, R3, and R4. By meeting these requirements, our algorithm effectively addresses the corresponding Challenges 1–3. No other baseline achieves comparable comprehensive performance across these dimensions. Thus, our algorithm stands out as a practical solution for urban low-altitude data collection scenarios.
Figure 5: The DCR, CR, and ECR about baselines and our algorithm when $N_{U^{\prime}}$ increases. The proposed algorithm consistently achieves the highest DCR with almost zero CR and lowest ECR, demonstrating superior efficiency and safety.
Figure 6: The DCR, CR, and SLR about baselines and our algorithm when $N_{GE}$ increases. Our algorithm maintains a stable and high DCR and CR of almost zero, while other baselines show declining performance and increased collisions. This highlights our algorithm’s robustness in complex environments.
Figure 7: The DCR, CR, and ECR about baselines and our algorithm when $N_{NFZ}$ increases. While other algorithms suffer from reduced data collection or increased collisions, our algorithm sustains near-perfect DCR and near zero CR, with competitive energy consumption.
Figure 8: The DCR, CR, and ECR about baselines and our algorithm when $N_{BZ}$ increases. The proposed algorithm preserves a high DCR and almost avoids collisions entirely, while other baselines experience noticeable performance degradation.
Figure 9: The DCR, CR, and RVR about baselines and our algorithm when $N_{RZ}$ increases. Our algorithm maintains full compliance (RVR 0), with high data collection efficiency and no collisions, while other baselines fails to adapt to regulatory constraints.
Figure 10: Average performance summary across all environmental configurations. Our algorithm achieves the highest DCR (99.49%), lowest CR (0%), and strong compliance (RVR 0% and ECR 76.95%), validating its overall effectiveness across diverse low-altitude scenarios.
7 Conclusion
In this paper, we have proposed a novel UAV trajectory planning algorithm tailored for low-altitude data collection in complex urban environments. Unlike prior works that focus on isolated objectives, our algorithm jointly addresses three key challenges: obstacle avoidance, regulatory compliance, and economically viable data collection. We have modeled the planning task as a partially observable Markov decision process and designed a hybrid optimization algorithm combining soft actor-critic algorithm with the large language model. Extensive experiments have demonstrated the effectiveness of our algorithm in improving data collection performance while reducing collisions, regulatory violations, and energy consumption.
Future work includes extending our algorithm to support multi-agent collaboration and real-world deployment in city-scale UAV networks. Moreover, integrating multi-modal sensor data and dynamically updating regulatory information will further enhance the generaliz ability and applicability of our algorithm.